# Recent Advances and Trends in Research Paper Recommender Systems: A Comprehensive Survey

### v1

@@ 1 Introduction
Global scientific publications are projected to continue growing significantly in the coming years. The overall market for scientific publishing is expected to grow at a Compound Annual Growth Rate (CAGR) of 3.3% from 2024 to 2030 [47]. For example, computer science publications are projected to reach 580,000 annually by 2025, reflecting a surge driven by advances in fields such as artificial intelligence and quantum computing [27]. In the case of the medical sector, it is expected to see a 5-7% annual increase in publications through 2025, driven by Artificial Intelligence (AI) and biotechnology. Meanwhile, the percentage of publications available through open access is forecasted to reach 45% by 2025, up from 38% in 2023 [5]. These trends reflect a broader, robust expansion in scientific research output across disciplines, particularly in computer science, fueled by technological advances and growing research investments.
@@
RP-RSs have evolved by adapting to new technologies, and although some of the underlying ideas and procedures remain, current systems are quite different from the first systems developed in the early 1990s. Deep Learning (DL) [56, 77, 70, 68, 32], and Large Languages Models (LLMs) [4, 52, 17, 55, 64] are more and more present in new proposals. In addition, Scientific Social Networks (SSNs) [67, 66, 21] also offer new opportunities that are undoubtedly being exploited. Despite the successful incorporation of all these advances, there is still significant room for improvement in both the evaluation approaches and the adoption of these systems in real world scenarios [57].
In this survey, an in-depth analysis of the current state of research on RP-RSs is provided, using as a starting point one of the most recent, complete and comprehensive surveys on the subject [35]. Since the work by [35] encompasses data up to October 2021, this new review extends the analysis to cover the period from November 2021 to December 2024, maintaining consistency by employing the same bibliographic review methodology. The main goal of this work is to answer the following research question:
@@ 2 Methodology
    The aim of this research question is to classify the works presented and to gain an understanding of how the systems generate the recommendations. In the literature, RP-RSs have traditionally been classified according to the taxonomy proposed by [14], even in literature reviews such as [9]. However, as stated by [35], this taxonomy has limitations to categorize RP-RSs that have been developed considering newer approaches and techniques. Therefore, this work classifies the RP-RSs considering the technology used and the purpose of the technique. To this end, RQ1 has been divided into the following questions:
@@
To answer these questions and to compare the current state of the research with previous surveys, the methodology implemented in [35], has been followed. The literature search has been conducted using the same digital libraries: ACM DL [^1] https://dl.acm.org/, DBLP [^2] https://dblp.uni-trier.de/, Google Scholar [^3] https://scholar.google.com/, and Springer [^4] https://link.springer.com/. The search query requested papers that contain the terms paper, article or publication in the title along with some form of recommendation, as in [35], and were published between 2021 and 2024. As the period considered in this work spans from November 2021 to December 2024, all the works published beyond that period were discarded.
@@ 3.1.1 What type of input do the RP-RSs use to generate the recommendations?
The third major kind of input is textual content (text), where a text that represents the interests of the users is used as input for the recommendation process. This approach has been followed by three works (4.76%). [62] uses paper abstracts as input. [56] takes a search query string as input. Finally, [11] uses different sections of the paper.
To a much lesser extent, combinations of the aforementioned types have also been found; [66] uses information of the research groups (group) in order to predict SSN group-paper ratings. [18] uses both the information of the paper and the user (user&paper) to predict the likelihood that user $u$ will cite paper $p$. [78] uses the information of authors and papers (user(a)&paper) to recommend papers by predicting the occurrence of co-authorship given a co-author ternary $(a_{1},p_{1},a_{2})$. Some works use author information and text (user(a)&text) to construct a query vector [26] or to search domain literature that reflects the characteristics and hot information of the domain [16]. [76] allows the users to specify their interests by introducing the authors, keywords, venues, organizations or funds (author||keyword||venue||organization||fund). Finally, [57] relies on user information along with text or papers and the publication year to represent the interests of the users (user&(text||papers)&publication year).
@@
paper [54, 65, 13, 43, 39, 23, 53, 58, 42, 80, 4, 22, 31, 64, 46, 52, 19, 38, 63] 30.16%
user [28, 67, 29, 33, 74, 15, 75, 25, 3, 81, 32, 24, 73, 44, 50, 60, 1, 61] 28.57%
user(a) [68, 17, 69, 77, 51, 30, 20, 40, 70, 21, 55, 45, 72, 49, 79, 71] 25.4%
text [11, 56, 62] 4.76%
user(a)&text [26, 16] 3.17%
user&paper [18] 1.59%
user(a)&paper [78] 1.59%
group [66] 1.59%
user&(text||papers)&publication year [57] 1.59%
author||keyword||venue||organization||fund [76] 1.59%
@@ 3.1.2 What kinds of information do the RP-RSs use to generate the recommendations?
Table LABEL:tab:PureContentInformation summarizes the papers that rely on the content explicitly included in the paper, along with information about the users, in the case of personalized recommendations [15], indicating the type of information used: title (Tit.), abstract (Abs.), keyword (Key), text (Txt.), author (Aut.), user (Us.), venue (Ven.), others (Oth.). The text category correspond to any textual content in the paper other than the title, abstract, authors and keywords.
Most of the works make use of the textual information contained in the papers. 68.75% of the papers use the title, 68.75% the abstract, and 18.75% use the keywords. 31.25% of the works use other kinds or unspecified type of text inputs. [13, 16] do not specify the kind of text used, while [53, 49] use the full text of the paper and [11] experiments with different parts of the text to examine the impact of text length (short, medium, and long texts are represented by abstract, abstract–introduction combination, and full text, respectively).
Author information is also widely used (37.5%). In particular, [49, 72, 16] use authors as input (user(a)) for the recommendation process. [64, 46] use the information of the authors to generate the representations of the papers. [76] uses authors as nodes in a graph-based approach. The use of venues (12.5%) is residual.
@@
[22, 52, 38]
[66, 63]
[13, 53]
[16, 49]
[11]
[46]
[15]
[64]
[72]
[76]
[62]
@@
Table LABEL:tab:RelationalInformation describes the RP-RSs which, besides the content of the paper, consider relational data to achieve their goal. Therefore, in addition to the information explicitly contained in the article, the use of relational data, such as interactions (Int.), citations (Cit.) and co-authors (Co-Aut.) is also analyzed. These relational data are used to improve the representation of papers and users, and therefore, generate better recommendations. For example, [45, 43, 30, 67] use the collaboration network of the authors to this end. [58] uses a hybrid similarity metric that combines text similarity and citation similarity to determine how close are the candidate paper and the POI. To achieve this goal, the citation network is used to calculate the similarity between the candidate papers and POI according to the articles that each of them refers to.
@@
The most commonly used relational data are citations (78.72%), probably because the information is easily accessible. 38.29% of the works rely on user information and 34.04% use interactions to generate the recommendations. In fact, most works offering personalized solutions based on users rely on interactions [74, 75, 73, 50, 1], such as browsing history or tagging, to extract user interests. In contrast, when the users are the authors themselves, citations serve as the primary source for extracting these interests [68, 17, 20, 69, 26, 79]. The co-author data is the least used of the three relational types (23.40%), and in most cases it is used together with citations [65, 31, 70, 45, 21, 71].
@@
[32, 3, 60]
[58, 23, 19]
[42, 54, 57]
[74, 75]
[24, 25]
[43, 45]
[77, 78]
[31, 21]
[81]
[17]
[30]
[44]
[18]
[20]
[65]
[33]
[55]
[39]
[70]
[80]
[4]
[26]
[61]
[40]
[56]
[71]
[1]
[51]
[67]
[69]
[73]
[79]
[50]
[68]
[28]
[29]
@@
Field of Study (FoS) has been used to generate paper embeddings [64, 65], capture researcher preferences in relation to papers and authors [4], and assist in clustering publications [57].
Topics are integrated into heterogeneous networks to obtain embeddings for papers and users [68, 69, 71]. Tags serve as nodes for meta-path creation and application of random walk algorithms [28, 29]. They are also fused with other information contained in the papers to extract meaningful features [74, 75, 1, 73].
Funding and organization data work similarly to tags, supporting meta-path creation and embedding generation in [76, 40]. Research groups are used in group-aware systems to create embeddings for papers, users, and groups [67, 66]. Claims are used as one part of an intra-document pair (e.g., paired with the title/abstract) in a patent dataset to form positive samples for contrastive learning [63]. Screenshots capture visual layout preferences, helping users select papers based on style [55]. The year indicates the novelty of the research [17], captures recency of papers, authors, and journals [80], and affects the decay of author interests, adjusting relevance in time-based heterogeneous graphs [51, 26]. Finally, named entities are used to create paper and user profiles [50].
@@
Year [17, 80, 51, 26, 40, 70, 54, 79] 12.70%
Tag [28, 29, 74, 75, 1, 73] 9.52%
Field of Study [64, 65, 4, 57] 6.35%
Topic [68, 69, 71] 4.76%
Organization [76, 40] 3.17%
Group [67, 66] 3.17%
Named entity [50] 1.59%
Fund [76] 1.59%
Screenshot [55] 1.59%
Claim [63] 1.59%

### v2

@@ 1 Introduction
Global scientific publications are projected to continue growing significantly in the coming years. The overall market for scientific publishing is expected to grow at a Compound Annual Growth Rate (CAGR) of 3.3% from 2024 to 2030 [OrbisResearch2023]. For example, computer science publications are projected to reach 580,000 annually by 2025, reflecting a surge driven by advances in fields such as artificial intelligence and quantum computing [PlutoInsights2025]. In the case of the medical sector, it is expected to see a 5-7% annual increase in publications through 2025, driven by Artificial Intelligence (AI) and biotechnology. Meanwhile, the percentage of publications available through open access is forecasted to reach 45% by 2025, up from 38% in 2023 [STM2024]. These trends reflect a broader, robust expansion in scientific research output across disciplines, particularly in computer science, fueled by technological advances and growing research investments.
@@
RP-RSs have evolved by adapting to new technologies, and although some of the underlying ideas and procedures remain, current systems are quite different from the first systems developed in the early 1990s. Deep Learning (DL) [IJ:DSM:Sivasankari:2023, IJ:ESWA:XIAO:2023, IJ:AS:Wang:2024, IC:DASFAA:Wang:2022, arXiv:KARIMI:2022], and Large Languages Models (LLMs) [IJ:Scientometrics:ALI:2022, IC:WIST:Roßner:2023, IJ:AI:GAO:2023, IC:ASSP:Shen:2024, IJ:Scientometrics:Thierry:2023] are more and more present in new proposals. In addition, Scientific Social Networks (SSNs) [IJ:ASC:Wang:2024, IJ:SMC:WANG:2022, IJ:mathematics:GUO:2024] also offer new opportunities that are undoubtedly being exploited. Despite the successful incorporation of all these advances, there is still significant room for improvement in both the evaluation approaches and the adoption of these systems in real world scenarios [IJ:KIS:STERGIOPOULOS:2024].
In this survey, an in-depth analysis of the current state of research on RP-RSs is provided, using as a starting point one of the most recent, complete and comprehensive surveys on the subject [IJ:DL:KREUTZ:2022]. Since the work by [IJ:DL:KREUTZ:2022] encompasses data up to October 2021, this new review extends the analysis to cover the period from November 2021 to December 2024, maintaining consistency by employing the same bibliographic review methodology. The main goal of this work is to answer the following research question:
@@ 2 Methodology
    The aim of this research question is to classify the works presented and to gain an understanding of how the systems generate the recommendations. In the literature, RP-RSs have traditionally been classified according to the taxonomy proposed by [burkeHybridRecommenderSystems2002], even in literature reviews such as [IJ:DL:BEEL:2015]. However, as stated by [IJ:DL:KREUTZ:2022], this taxonomy has limitations to categorize RP-RSs that have been developed considering newer approaches and techniques. Therefore, this work classifies the RP-RSs considering the technology used and the purpose of the technique. To this end, RQ1 has been divided into the following questions:
@@
To answer these questions and to compare the current state of the research with previous surveys, the methodology implemented in [IJ:DL:KREUTZ:2022], has been followed. The literature search has been conducted using the same digital libraries: ACM DL [^1] https://dl.acm.org/, DBLP [^2] https://dblp.uni-trier.de/, Google Scholar [^3] https://scholar.google.com/, and Springer [^4] https://link.springer.com/. The search query requested papers that contain the terms paper, article or publication in the title along with some form of recommendation, as in [IJ:DL:KREUTZ:2022], and were published between 2021 and 2024. As the period considered in this work spans from November 2021 to December 2024, all the works published beyond that period were discarded.
@@ 3.1.1 What type of input do the RP-RSs use to generate the recommendations?
The third major kind of input is textual content (text), where a text that represents the interests of the users is used as input for the recommendation process. This approach has been followed by three works (4.76%). [EC:ECIR:Takahashi:2022] uses paper abstracts as input. [IJ:DSM:Sivasankari:2023] takes a search query string as input. Finally, [IJ:ELECTRONICS:Blazevic:2023] uses different sections of the paper.
To a much lesser extent, combinations of the aforementioned types have also been found; [IJ:SMC:WANG:2022] uses information of the research groups (group) in order to predict SSN group-paper ratings. [IC:MLBDBI:GUAN:2022] uses both the information of the paper and the user (user&paper) to predict the likelihood that user $u$ will cite paper $p$. [IJ:KBS:Xiao:2023] uses the information of authors and papers (user(a)&paper) to recommend papers by predicting the occurrence of co-authorship given a co-author ternary $(a_{1},p_{1},a_{2})$. Some works use author information and text (user(a)&text) to construct a query vector [IJ:Scientometrics:Huang:2024] or to search domain literature that reflects the characteristics and hot information of the domain [IJ:AS:Chen:2023]. [IJ:AI:Xiao:2023] allows the users to specify their interests by introducing the authors, keywords, venues, organizations or funds (author||keyword||venue||organization||fund). Finally, [IJ:KIS:STERGIOPOULOS:2024] relies on user information along with text or papers and the publication year to represent the interests of the users (user&(text||papers)&publication year).
@@
paper [ARXIV:Shen:2024, IJ:AI:Thierry:2023, IJ:EIT:Smail:2023, IJ:ESWA:MEI:2022, IJ:Heliyon:LI:2024, IJ:IEEE:HADHIATMA:2023, IJ:IEEE:SARWAR:2021, IJ:IEEE:STERLING:2022, IJ:IPM:Long:2024, IJ:JOI:Yadav:2022, IJ:Scientometrics:ALI:2022, IJ:Scientometrics:GUNDOGAN:2022, IJ:Scientometrics:Kanwal:2024, IJ:Scientometrics:Thierry:2023, IJ:SECS:Hamisu:2024, IC:WIST:Roßner:2023, IJ:SHTI:GUO:2022, IJ:AMEC:KUS:2023, IJ:Scientometrics:TAN:2023] 30.16%
user [IC:IKT:JAFARI:2021, IJ:ASC:Wang:2024, IJ:AUT:Jafari:2023, IJ:DATABASE:Kart:2022, IJ:ESWA:XI:2023, IJ:IS:Chaudhuri:2022, IJ:SC:XI:2024, J:EIJ:Huang:2024, IJ:CDS:AHMEDI:2022, IC:KSEM:YU:2022, arXiv:KARIMI:2022, IJ:JS:Huang:2023, IS:CSET:WU:2022, J:JIS:MUNGEN:2022, IJ:IS:PINEDO:2024, IC:ICITACEE:Switrayana:2022, ARXIV:Mohamed:2022, PP:Mahesh:2023] 28.57%
user(a) [IC:DASFAA:Wang:2022, IJ:AI:GAO:2023, IJ:ESWA:Wang:2023, IJ:ESWA:XIAO:2023, IJ:IEEE:REN:2022, IJ:Scientometrics:Jiang:2023, IJ:TIIS:GUO:2024, CF:CSCW:LI:2022, IJ:AS:Wang:2024, IJ:mathematics:GUO:2024, IC:ASSP:Shen:2024, IC:EITCE:NIU:2023, IJ:JES:LI:2024, AC:ACIIDS:PAN:2024, IC:ICDE:XIE:2022, IC:SEAI:Wang:2024] 25.4%
text [IJ:ELECTRONICS:Blazevic:2023, IJ:DSM:Sivasankari:2023, EC:ECIR:Takahashi:2022] 4.76%
user(a)&text [IJ:Scientometrics:Huang:2024, IJ:AS:Chen:2023] 3.17%
user&paper [IC:MLBDBI:GUAN:2022] 1.59%
user(a)&paper [IJ:KBS:Xiao:2023] 1.59%
group [IJ:SMC:WANG:2022] 1.59%
user&(text||papers)&publication year [IJ:KIS:STERGIOPOULOS:2024] 1.59%
author||keyword||venue||organization||fund [IJ:AI:Xiao:2023] 1.59%
@@ 3.1.2 What kinds of information do the RP-RSs use to generate the recommendations?
Table LABEL:tab:PureContentInformation summarizes the papers that rely on the content explicitly included in the paper, along with information about the users, in the case of personalized recommendations [IJ:IS:Chaudhuri:2022], indicating the type of information used: title (Tit.), abstract (Abs.), keyword (Key), text (Txt.), author (Aut.), user (Us.), venue (Ven.), others (Oth.). The text category correspond to any textual content in the paper other than the title, abstract, authors and keywords.
Most of the works make use of the textual information contained in the papers. 68.75% of the papers use the title, 68.75% the abstract, and 18.75% use the keywords. 31.25% of the works use other kinds or unspecified type of text inputs. [IJ:EIT:Smail:2023, IJ:AS:Chen:2023] do not specify the kind of text used, while [IJ:IEEE:SARWAR:2021, AC:ACIIDS:PAN:2024] use the full text of the paper and [IJ:ELECTRONICS:Blazevic:2023] experiments with different parts of the text to examine the impact of text length (short, medium, and long texts are represented by abstract, abstract–introduction combination, and full text, respectively).
Author information is also widely used (37.5%). In particular, [AC:ACIIDS:PAN:2024, IJ:JES:LI:2024, IJ:AS:Chen:2023] use authors as input (user(a)) for the recommendation process. [IJ:Scientometrics:Thierry:2023, IJ:SECS:Hamisu:2024] use the information of the authors to generate the representations of the papers. [IJ:AI:Xiao:2023] uses authors as nodes in a graph-based approach. The use of venues (12.5%) is residual.
@@
[IJ:Scientometrics:GUNDOGAN:2022, IC:WIST:Roßner:2023, IJ:AMEC:KUS:2023]
[IJ:SMC:WANG:2022, IJ:Scientometrics:TAN:2023]
[IJ:EIT:Smail:2023, IJ:IEEE:SARWAR:2021]
[IJ:AS:Chen:2023, AC:ACIIDS:PAN:2024]
[IJ:ELECTRONICS:Blazevic:2023]
[IJ:SECS:Hamisu:2024]
[IJ:IS:Chaudhuri:2022]
[IJ:Scientometrics:Thierry:2023]
[IJ:JES:LI:2024]
[IJ:AI:Xiao:2023]
[EC:ECIR:Takahashi:2022]
@@
Table LABEL:tab:RelationalInformation describes the RP-RSs which, besides the content of the paper, consider relational data to achieve their goal. Therefore, in addition to the information explicitly contained in the article, the use of relational data, such as interactions (Int.), citations (Cit.) and co-authors (Co-Aut.) is also analyzed. These relational data are used to improve the representation of papers and users, and therefore, generate better recommendations. For example, [IC:EITCE:NIU:2023, IJ:ESWA:MEI:2022, IJ:Scientometrics:Jiang:2023, IJ:ASC:Wang:2024] use the collaboration network of the authors to this end. [IJ:IEEE:STERLING:2022] uses a hybrid similarity metric that combines text similarity and citation similarity to determine how close are the candidate paper and the POI. To achieve this goal, the citation network is used to calculate the similarity between the candidate papers and POI according to the articles that each of them refers to.
@@
The most commonly used relational data are citations (78.72%), probably because the information is easily accessible. 38.29% of the works rely on user information and 34.04% use interactions to generate the recommendations. In fact, most works offering personalized solutions based on users rely on interactions [IJ:ESWA:XI:2023, IJ:SC:XI:2024, IS:CSET:WU:2022, IJ:IS:PINEDO:2024, ARXIV:Mohamed:2022], such as browsing history or tagging, to extract user interests. In contrast, when the users are the authors themselves, citations serve as the primary source for extracting these interests [IC:DASFAA:Wang:2022, IJ:AI:GAO:2023, IJ:TIIS:GUO:2024, IJ:ESWA:Wang:2023, IJ:Scientometrics:Huang:2024, IC:ICDE:XIE:2022]. The co-author data is the least used of the three relational types (23.40%), and in most cases it is used together with citations [IJ:AI:Thierry:2023, IJ:Scientometrics:Kanwal:2024, IJ:AS:Wang:2024, IC:EITCE:NIU:2023, IJ:mathematics:GUO:2024, IC:SEAI:Wang:2024].
@@
[arXiv:KARIMI:2022, IJ:CDS:AHMEDI:2022, IC:ICITACEE:Switrayana:2022]
[IJ:IEEE:STERLING:2022, IJ:IEEE:HADHIATMA:2023, IJ:SHTI:GUO:2022]
[IJ:IPM:Long:2024, ARXIV:Shen:2024, IJ:KIS:STERGIOPOULOS:2024]
[IJ:ESWA:XI:2023, IJ:SC:XI:2024]
[IJ:JS:Huang:2023, J:EIJ:Huang:2024]
[IJ:ESWA:MEI:2022, IC:EITCE:NIU:2023]
[IJ:ESWA:XIAO:2023, IJ:KBS:Xiao:2023]
[IJ:Scientometrics:Kanwal:2024, IJ:mathematics:GUO:2024]
[IC:KSEM:YU:2022]
[IJ:AI:GAO:2023]
[IJ:Scientometrics:Jiang:2023]
[J:JIS:MUNGEN:2022]
[IC:MLBDBI:GUAN:2022]
[IJ:TIIS:GUO:2024]
[IJ:AI:Thierry:2023]
[IJ:DATABASE:Kart:2022]
[IC:ASSP:Shen:2024]
[IJ:Heliyon:LI:2024]
[IJ:AS:Wang:2024]
[IJ:JOI:Yadav:2022]
[IJ:Scientometrics:ALI:2022]
[IJ:Scientometrics:Huang:2024]
[$P$:Mahesh:2023]
[CF:CSCW:LI:2022]
[IJ:DSM:Sivasankari:2023]
[IC:SEAI:Wang:2024]
[ARXIV:Mohamed:2022]
[IJ:IEEE:REN:2022]
[IJ:ASC:Wang:2024]
[IJ:ESWA:Wang:2023]
[IS:CSET:WU:2022]
[IC:ICDE:XIE:2022]
[IJ:IS:PINEDO:2024]
[IC:DASFAA:Wang:2022]
[IC:IKT:JAFARI:2021]
[IJ:AUT:Jafari:2023]
@@
Field of Study (FoS) has been used to generate paper embeddings [IJ:Scientometrics:Thierry:2023, IJ:AI:Thierry:2023], capture researcher preferences in relation to papers and authors [IJ:Scientometrics:ALI:2022], and assist in clustering publications [IJ:KIS:STERGIOPOULOS:2024].
Topics are integrated into heterogeneous networks to obtain embeddings for papers and users [IC:DASFAA:Wang:2022, IJ:ESWA:Wang:2023, IC:SEAI:Wang:2024]. Tags serve as nodes for meta-path creation and application of random walk algorithms [IC:IKT:JAFARI:2021, IJ:AUT:Jafari:2023]. They are also fused with other information contained in the papers to extract meaningful features [IJ:ESWA:XI:2023, IJ:SC:XI:2024, ARXIV:Mohamed:2022, IS:CSET:WU:2022].
Funding and organization data work similarly to tags, supporting meta-path creation and embedding generation in [IJ:AI:Xiao:2023, CF:CSCW:LI:2022]. Research groups are used in group-aware systems to create embeddings for papers, users, and groups [IJ:ASC:Wang:2024, IJ:SMC:WANG:2022]. Claims are used as one part of an intra-document pair (e.g., paired with the title/abstract) in a patent dataset to form positive samples for contrastive learning [IJ:Scientometrics:TAN:2023]. Screenshots capture visual layout preferences, helping users select papers based on style [IC:ASSP:Shen:2024]. The year indicates the novelty of the research [IJ:AI:GAO:2023], captures recency of papers, authors, and journals [IJ:JOI:Yadav:2022], and affects the decay of author interests, adjusting relevance in time-based heterogeneous graphs [IJ:IEEE:REN:2022, IJ:Scientometrics:Huang:2024]. Finally, named entities are used to create paper and user profiles [IJ:IS:PINEDO:2024].
@@
Year [IJ:AI:GAO:2023, IJ:JOI:Yadav:2022, IJ:IEEE:REN:2022, IJ:Scientometrics:Huang:2024, CF:CSCW:LI:2022, IJ:AS:Wang:2024, ARXIV:Shen:2024, IC:ICDE:XIE:2022] 12.70%
Tag [IC:IKT:JAFARI:2021, IJ:AUT:Jafari:2023, IJ:ESWA:XI:2023, IJ:SC:XI:2024, ARXIV:Mohamed:2022, IS:CSET:WU:2022] 9.52%
Field of Study [IJ:Scientometrics:Thierry:2023, IJ:AI:Thierry:2023, IJ:Scientometrics:ALI:2022, IJ:KIS:STERGIOPOULOS:2024] 6.35%
Topic [IC:DASFAA:Wang:2022, IJ:ESWA:Wang:2023, IC:SEAI:Wang:2024] 4.76%
Organization [IJ:AI:Xiao:2023, CF:CSCW:LI:2022] 3.17%
Group [IJ:ASC:Wang:2024, IJ:SMC:WANG:2022] 3.17%
Named entity [IJ:IS:PINEDO:2024] 1.59%
Fund [IJ:AI:Xiao:2023] 1.59%
Screenshot [IC:ASSP:Shen:2024] 1.59%
Claim [IJ:Scientometrics:TAN:2023] 1.59%